\begin{thebibliography}{CMDGF21}

\bibitem[Bar23]{Barajas2023}
Paul Barajas.
\newblock High-order derivations of the hasse-schmidt algebra.
\newblock {\em arXiv:2305.04537}, 2023.

\bibitem[BCCD25]{BCCD}
Wágner Badilla-Céspedes, Abel Castorena, and Daniel Duarte.
\newblock Higher jacobian matrix of weighted homogeneous polynomials and derivation algebras.
\newblock {\em Journal of Singularities}, 28:197--206, 2025.

\bibitem[BD20]{BD2020}
Paul Barajas and Daniel Duarte.
\newblock On the module of differentials of order {$n$} of hypersurfaces.
\newblock {\em J. Pure Appl. Algebra}, 224(2):536--550, 2020.

\bibitem[BJNB19]{BJNB2019}
Holger Brenner, Jack Jeffries, and Luis N\'{u}{\~n}ez-Betancourt.
\newblock Quantifying singularities with differential operators.
\newblock {\em Adv. Math.}, 358:106843, 89, 2019.

\bibitem[CCYZ20]{CCYZ2020}
Bingyi Chen, Hao Chen, Stephen S.-T. Yau, and Huaiqing Zuo.
\newblock The nonexistence of negative weight derivations on positive dimensional isolated singularities: generalized {W}ahl conjecture.
\newblock {\em J. Differential Geom.}, 115(2):195--224, 2020.

\bibitem[Che99]{H1999}
Hao Chen.
\newblock Nonexistence of negative weight derivations on graded {A}rtin algebras: a conjecture of {H}alperin.
\newblock {\em J. Algebra}, 216(1):1--12, 1999.

\bibitem[CMDGF21]{MR4229623}
Enrique Ch\'{a}vez-Mart\'{\i}nez, Daniel Duarte, and Arturo Giles~Flores.
\newblock A higher-order tangent map and a conjecture on the higher {N}ash blowup of curves.
\newblock {\em Math. Z.}, 297(3-4):1767--1791, 2021.

\bibitem[CXY95]{CXY1995}
Hao Chen, Yi-Jing Xu, and Stephen S.-T. Yau.
\newblock Nonexistence of negative weight derivation of moduli algebras of weighted homogeneous singularities.
\newblock {\em J. Algebra}, 172(2):243--254, 1995.

\bibitem[CYZ19]{CYZ2019}
Hao Chen, Stephen S.-T. Yau, and Huaiqing Zuo.
\newblock Non-existence of negative weight derivations on positively graded {A}rtinian algebras.
\newblock {\em Trans. Amer. Math. Soc.}, 372(4):2493--2535, 2019.

\bibitem[dAD21]{AD2021}
Hern\'{a}n de~Alba and Daniel Duarte.
\newblock On the {$k$}-torsion of the module of differentials of order {$n$} of hypersurfaces.
\newblock {\em J. Pure Appl. Algebra}, 225(8):Paper No. 106646, 7, 2021.

\bibitem[Dua17]{Duarte2017}
Daniel Duarte.
\newblock Computational aspects of the higher {N}ash blowup of hypersurfaces.
\newblock {\em J. Algebra}, 477:211--230, 2017.

\bibitem[HMYZ23]{HMYZ2023}
Naveed Hussain, Guorui Ma, Stephen S.-T. Yau, and Huaiqing Zuo.
\newblock Higher {N}ash blow-up local algebras of singularities and its derivation {L}ie algebras.
\newblock {\em J. Algebra}, 618:165--194, 2023.

\bibitem[LDS24]{LDS2023}
Yann Le~Dr\'eau and Julien Sebag.
\newblock Arc scheme and higher differential forms.
\newblock {\em Osaka J. Math.}, 61(3):381--390, 2024.

\bibitem[LY25]{LeYasuda2025}
Quy~Thuong L\^e and Takehiko Yasuda.
\newblock Higher {J}acobian ideals, contact equivalence and motivic zeta functions.
\newblock {\em Rev. Mat. Iberoam.}, 41(3):1057--1080, 2025.

\bibitem[MY82]{MR0674404}
John~N. Mather and Stephen S.~T. Yau.
\newblock Classification of isolated hypersurface singularities by their moduli algebras.
\newblock {\em Invent. Math.}, 69(2):243--251, 1982.

\bibitem[Sai71]{Saito1971}
Kyoji Saito.
\newblock Quasihomogene isolierte {S}ingularit\"{a}ten von {H}yperfl\"{a}chen.
\newblock {\em Invent. Math.}, 14:123--142, 1971.

\bibitem[XY96]{XY1996}
Yi-Jing Xu and Stephen S.-T. Yau.
\newblock Micro-local characterization of quasi-homogeneous singularities.
\newblock {\em Amer. J. Math.}, 118(2):389--399, 1996.

\bibitem[Yau91]{Y1991}
Stephen S.-T. Yau.
\newblock Solvability of {L}ie algebras arising from isolated singularities and nonisolatedness of singularities defined by {${\rm sl}(2,{\bf C})$} invariant polynomials.
\newblock {\em Amer. J. Math.}, 113(5):773--778, 1991.

\bibitem[Ye25]{FeiYe}
Fei Ye.
\newblock On the structure of second jacobian ideals.
\newblock {\em Beitr Algebra Geom}, 2025.

\bibitem[YZ16]{MR3490075}
Stephen S.-T. Yau and Huaiqing Zuo.
\newblock Derivations of the moduli algebras of weighted homogeneous hypersurface singularities.
\newblock {\em J. Algebra}, 457:18--25, 2016.

\end{thebibliography}

\documentclass[12pt]{amsart}

 \usepackage[pagewise]{lineno}
 
\usepackage[colorlinks=true,pagebackref,hyperindex,citecolor=blue,linkcolor=red]{hyperref}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{setspace}
\usepackage{moreverb}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{amssymb}
\usepackage{tikz}
\usepackage{color}
\usepackage[all]{xy}
\usepackage{xfrac}
\usepackage[top=1in, bottom=1in, left=1in, right=1in]{geometry}
\usepackage{mathrsfs}

\usetikzlibrary{shapes}

\newtheorem{theoremx}{Theorem}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{example}[theorem]{Example}

\newtheorem{conjecture}[theorem]{Conjecture}

\newtheorem{remark}[theorem]{Remark}
\numberwithin{equation}{subsection}

\newcommand{\NN}{\mathbb{N}}
\newcommand{\CC}{\mathbb{C}}

\newcommand{\cM}{\mathcal{T}}
\newcommand{\cL}{\mathcal{L}}

\newcommand{\cJ}{\mathcal{J}}
\newcommand{\cT}{\mathcal{T}}

\newcommand{\pT}{\CC[x_1,\ldots,x_s]/\langle f,\cJ_n(f)\rangle}

\newcommand{\Der}{\operatorname{Der}}
\newcommand{\Jac}{\operatorname{Jac}}
\newcommand{\conv}{\CC\{x_1,\ldots,x_s\}}

\newcommand{\pol}{\CC[x_1,\ldots,x_s]}

\newcommand{\dw}{\deg_w}

\begin{document}
\newcommand{\tens}{\otimes}
\newcommand{\hhtest}[1]{\tau ( #1 )}
\renewcommand{\hom}[3]{\operatorname{Hom}_{#1} ( #2, #3 )}

\allowdisplaybreaks

\title[Non-existence of negative derivations]{Non-existence of negative derivations on the higher Nash blowup local algebra}

\author[W. Badilla-C\'espedes]{Wágner Badilla-Céspedes}
\address{Centro de Ciencias Matemáticas, UNAM, Campus Morelia, Morelia, Michoacán, México.}
\email{wagner@matmor.unam.mx }

\author[A. Castorena]{Abel Castorena}

\email{abel@matmor.unam.mx}

\author[D. Duarte]{Daniel Duarte}

\email{adduarte@matmor.unam.mx}
\author[L. Núñez-Betancourt]{Luis Núñez-Betancourt}
\address{Centro de Investigación en Matemáticas, Guanajuato, Gto., México.}
\email{luisnub@cimat.mx}

\subjclass[2020]{14B05, 32S05, 13N15.}
\keywords{Higher Nash blowup local algebras, higher-order Jacobian matrix, weighted homogeneous isolated hypersurface singularities}

\begin{abstract}
Let $f\in\pol$ be a weighted homogeneous polynomial having an isolated singularity and $\cT_n(f)$ be its higher Nash blowup local algebra. We show that $\cT_n(f)$ does not admit negative weighted derivations for $n\geq2$. This answers affirmatively a conjecture of Hussain-Ma-Yau-Zuo.
\end{abstract}

\maketitle

\section*{Introduction}

The  Jacobian matrix of order $n$ is a higher-order version of the classical Jacobian matrix. It was introduced as a tool for computing the higher Nash blowup of a hypersurface \cite{Duarte2017}. Higher-order Jacobian matrices were later rediscovered and further developed by several authors \cite{BJNB2019,BD2020,MR4229623}. Ever since, this matrix has seen a wide variety of applications in singularity theory \cite{Duarte2017,BJNB2019,BD2020,MR4229623,AD2021,Barajas2023,HMYZ2023,LDS2023,LeYasuda2025,FeiYe}.

Let $f\in\conv$, where $\conv$ denotes the ring of convergent power series at the origin. Let $\Jac_n(f)$ be the higher Jacobian matrix of $f$ and $\cJ_n(f)$ the ideal generated by the maximal minors of $\Jac_n(f)$ (see Definition \ref{Jacn}). Hussain-Ma-Yau-Zuo defined the higher Nash blowup local algebra as the quotient $\cM_n(f)=\conv/\langle f,\cJ_n(f) \rangle$ \cite{HMYZ2023}. For $n=1$, this is the classical Tjurina algebra. Moreover, the authors also introduced the  algebra of derivations $\cL_n(f)=\Der_{\CC}(\cM_n(f))$. For $n=1$, this is known as the Yau algebra of $f$. The Tjurina and Yau algebras are fundamental objects in the study of hypersurface singularities \cite{MR0674404,Y1991}.

Three conjectures were proposed regarding $\cT_n(f)$: invariance under contact equivalence, non-existence of negative weighted derivations on $\cT_n(f)$ whenever $f$ is weighted homogeneous, and bounds for the dimension of $\cL_n(f)$ \cite{HMYZ2023}. 


The conjecture regarding the invariance of $\cT_n(f)$ under contact equivalence was proved by L\^{e}-Yasuda \cite{LeYasuda2025}. In this paper we prove the conjecture on the non-existence of negative weighted derivations on $\cT_n(f)$. Let us discuss this conjecture in more detail.
 
It is known that $\pT$ is a graded algebra whenever $f$ is a weighted homogeneous polynomial \cite[Theorem 1.7]{BCCD}. In this case, $\cL_n(f)$ is also graded. There is a general conjecture on algebras of derivations of local Artinian graded algebras, known as Halperin conjecture, stating the non-existence of negative weight derivations over those algebras \cite{H1999}. Inspired by Halperin's question, Hussain-Ma-Yau-Zuo conjectured that there are no negative weighted derivations on $\pT$. The goal of this paper is to prove this conjecture.

\begin{theoremx}\label{theorem-main}
Let $f\in \CC[x_1,\ldots,x_s]$ be a weighted homogeneous polynomial with respect to $w\in\NN^s$ and degree $d$ that defines an isolated hypersurface singularity. Assume that $d\geq2w_1\geq\cdots\geq 2w_s>0$. Then $\pT$ does not admit negative weighted derivations for all $n\geq2$.
\end{theoremx}

\section*{Acknowledgements}
The first author is supported by UNAM Posdoctoral Program (POSDOC). The second author is supported by project IN100723, “Curvas, Sistemas lineales en superficies proyectivas y fibrados vectoriales” from DGAPA, UNAM. The third author is supported by SECIHTI project CF-2023-G-33 and PAPIIT grant IN117523. 
The fourth author was  supported by SECIHTI Grants CBF 2023-2024-224 and CF-2023-G-33.

\section{Higher-order Jacobian matrix of a weighted homogeneous polynomial}

\begin{definition}[{\cite{Duarte2017,BJNB2019,BD2020}}] \label{Jacn}
Let $f\in\conv$. Denote
$$\Jac_n(f):=\left(\frac{1}{(\alpha-\beta)!}\frac{\partial^{\alpha-\beta}(f)}{\partial x^{\alpha-\beta}}\right)_{\substack{\beta\in \{\beta\in\NN^s|0\leq|\beta|\leq n-1\} \\ \alpha\in \{\alpha\in\NN^s|1\leq|\alpha|\leq n\}}},$$
where we define $\displaystyle{\frac{1}{(\alpha-\beta)!}\frac{\partial^{\alpha-\beta}(f)}{\partial x^{\alpha-\beta}}=0},$ whenever $\alpha_i<\beta_i$ for some $i$ or if $\beta=\alpha$. It is a $M\times(N-1)$-matrix, where $M=\binom{n+s-1}{s}$ and $N=\binom{n+s}{s}$. We call $\Jac_n(f)$ the Jacobian matrix of order $n$ of $f$. 

Moreover, denote as $\cJ_n(f)$ the ideal generated by all maximal minors of $\Jac_n(f)$. Notice that $\Jac_1(f)$ is the usual Jacobian matrix and $\cJ_1(f)$ is the usual Jacobian ideal of $f$.
\end{definition}

\begin{example}\label{j2}
Let $f\in\CC[x_1,x_2]$. Denote as $f_i=\frac{\partial f}{\partial x_i}$ and $f_{ij}=\frac{\partial^2 f}{\partial x_ix_j}$. Then 
\[\Jac_2(f)=
\left( 
\begin{array}{ccccc}
             f_1& f_2 & \frac{1}{2}f_{11} & f_{12} & \frac{1}{2}f_{22} \\
             0 & 0 & f_1 & f_2 & 0 \\
             0 & 0 & 0 & f_1 & f_2\\
\end{array} 
\right).\] 
Moreover,
$$\cJ_2(f)=\langle f_1^3,f_1^2f_2,f_1f_2^2,f_2^3,\frac{1}{2}f_1^2f_{22}-f_1f_2f_{12}+\frac{1}{2}f_2^2f_{11}\rangle\subset\CC[x_1,x_2].$$
\end{example}

Recall that a polynomial $f=\sum_{\alpha}c_{\alpha}x^{\alpha}\in\pol$ is called weighted homogeneous of weighted degree $d$ with respect to $w\in\NN_{\geq1}^s$, if $\alpha\cdot w=d$ for all $c_{\alpha}\neq0$. We denote the weighted degree of $f$ as $\dw(f)$. 


\begin{theorem}[{\cite[Theorem 1.7]{BCCD}}]\label{theorem:Tjurina-graded}
Let $f$ be a weighted homogeneous polynomial in $\CC[x_1,\ldots,x_s]$. Then $\cJ_n(f)$ is a weighted homogeneous ideal for every $n \in \NN$.   
\end{theorem}

Hussain-Ma-Yau-Zuo used the ideal $\cJ_n(f)$ to introduce the following generalization of the Tjurina algebra of a hypersurface.

\begin{definition}[{\cite[Definition 1.3]{HMYZ2023}}] \label{Tjurina}
Let $f\in\conv$. Denote
$$\cT_n(f)=\conv/\langle f,\cJ_n(f) \rangle.$$
$\cT_n(f)$ is called the higher Nash blowup local algebra of $f$. Moreover, denote $\cL_n(f)=\Der_{\CC}(\cT_n(f))$, i.e., the Lie algebra of derivations of $\cT_n(f)$. Notice that $\cT_1(f)$ and $\cL_1(f)$ are the Tjurina algebra and the Yau algebra of $f$, respectively.
\end{definition}

\begin{remark}
There are two versions of the higher-order Jacobian matrix in the literature. The original one has $f$ in the $(\beta,\beta)$-entry of $\Jac_n(f)$ \cite{Duarte2017}, whereas the definition given in \cite{BJNB2019,BD2020} has 0 in the $(\beta,\beta)$-entry. Both versions coincide modulo $f$. In particular, the definition of $\cT_n(f)$ is the same regardless of the version that is adopted.
\end{remark}

In view of Theorem \ref{theorem:Tjurina-graded}, the algebra of derivations of $\pT$ is naturally graded. Hussain-Ma-Yau-Zuo proposed the following conjecture, which they verified for $s=n=2$ \cite[Theorem B]{HMYZ2023}. Later, the conjecture was verified for $s=2$ and $n\geq2$ \cite[Theorem 2.6]{BCCD}.

\begin{conjecture}[{\cite[Conjecture 1.7]{HMYZ2023}}]\label{conjecture}
Let $f\in\CC[x_1,\ldots,x_s]$ be a weighted homogeneous polynomial defining a hypersurface with  isolated singularity. Assume that $\dw(f)\geq 2w_1\geq2w_2\geq\cdots\geq2w_s>0$. Then $\pT$ has no negative weighted derivations, for $n\geq2$.
\end{conjecture}

\begin{remark}
Recall that $\CC[x_1,\ldots,x_s]/\langle f, \mathcal{J}_1(f)\rangle$ coincides with the classical Tjurina algebra. The analogous statement of the previous conjecture for the Tjurina algebra is a conjecture formulated by Yau \cite{CXY1995}. To the best of our knowledge, Yau's conjecture is still open, although many cases have already been resolved \cite{CXY1995,XY1996,CYZ2019,CCYZ2020}.
\end{remark}

\section{Proof of Theorem \ref{theorem-main}}

\begin{proposition}\label{leastdeg}
Let $f\in\pol$ be a weighted homogeneous polynomial with respect to $w\in\NN_{\geq1}^s$. Assume that $\dw(f)\geq 2w_1\geq2w_2\geq\cdots\geq2w_s>0$. Then
\begin{equation*}
\dw(f)\leq \dw(g),
\end{equation*}
where g denotes an arbitrary maximal minor of $\Jac_n(f)$, for $n\geq2$.
\end{proposition}
\begin{proof}
This can be checked directly for $s=n=2$ using the generators from Example \ref{j2}. Hence we can assume $s\geq3$ and $n\geq2$, or $s=2$ and $n\geq3$. In both cases, it is known that $\cJ_n(f)\subseteq\cJ_1(f)^3$ \cite[Proposition 2.20]{LeYasuda2025}. Let $d=\deg_w(f)$.

Denote $f_i=\frac{\partial f}{\partial x_i}$. Let $h=f_1^{\alpha_1}\cdots f_s^{\alpha_s}$ be a generator of $\cJ_1(f)^3$, where $\alpha_1+\cdots+\alpha_s=3$. Using the assumption $d\geq 2w_1\geq\cdots\geq2w_s>0$ we obtain
\begin{align*}
\deg_w(h)&=\alpha_1(d-w_1)+\cdots+\alpha_s(d-w_s)\\
&\geq \alpha_1(d-w_1)+\cdots+\alpha_s(d-w_1)\\
&=(\alpha_1+\cdots+\alpha_s)(d-w_1)\\
&=3(d-w_1)=d+2d-3w_1\\
&\geq d+4w_1-3w_1=d+w_1\geq d.
\end{align*}
Since $\cJ_1(f)^3$ is weighted homogeneous, the previous computation shows that any non-zero weighted element of $\cJ_1(f)^3$ has weighted degree greater or equal than $d$. In particular, this is true for the generators of $\cJ_n(f)$.
\end{proof}


\begin{proof}[Proof of Theorem \ref{theorem-main}]
Since $f$ has an isolated singularity and is weighted homogeneous, the ideal generated by the partial derivatives $f_i=\frac{\partial f}{\partial x_i}$ is zero-dimensional. Then
$f_1,\ldots,f_s$ form a weighted homogeneous regular sequence of degrees $d-w_1,\ldots,d-w_s$, respectively. 

Let $\delta\in \Der_\CC (\pT)$ be a weighted homogeneous derivation of negative degree. Then $\delta$ has a weighted homogeneous lifting
$\tilde{\delta}\in \Der_\CC(\pol)$ of the same degree such that $\tilde{\delta}(\langle f,\cJ_n(f) \rangle)\subseteq \langle f,\cJ_n(f) \rangle$. Since $\tilde{\delta}$ has negative weighted degree, Proposition \ref{leastdeg} implies $\tilde{\delta}(f)=0.$
Let $r_1,\ldots,r_s\in \pol$ be such that
$\tilde{\delta}=r_1\frac{\partial}{\partial x_1}+\ldots+r_s\frac{\partial}{\partial x_s}$. Then, 
$$0=\tilde{\delta}(f)=r_1 f_1+\ldots+r_s f_s.$$
Since $f_1,\ldots,f_s$ is a regular sequence, we have that 
$$
\begin{pmatrix}
r_1\\
r_2\\
\vdots\\
r_s
\end{pmatrix}\in \langle f_ie_j-f_je_i\rangle_{i\neq j},
$$
where $e_i$ is the $i$-th vector in the canonical basis of $(\pol)^s$.
Hence, 
$$
\tilde{\delta}=r_1\frac{\partial}{\partial x_1}+\dots+r_s\frac{\partial}{\partial x_s}\in
\left\langle 
f_i \frac{\partial}{\partial x_j}-f_j\frac{\partial}{\partial x_i}\right\rangle_{i\neq j}.
$$
Then,
\begin{align*}
\dw(\delta)=\dw(\tilde{\delta})\geq &\min\left\{\dw \left( f_i \frac{\partial}{\partial x_j}-f_j\frac{\partial}{\partial x_i}\right)\right\}_{i\neq j}\\
=&\min\left\{ \dw(f)-w_i-w_j\right\}_{i\neq j}\geq 0.    
\end{align*}
This contradicts the fact that $\delta$ has negative degree.
\end{proof}

\begin{remark}
The assumption $\dw(f)\geq 2w_1\geq2w_2\geq\cdots\geq2w_s>0$ holds, up to a change of coordinates, whenever $f\in\langle x_1,\ldots,x_s\rangle^3$ \cite[Theorem 2.1]{MR3490075}, \cite{Saito1971}.
\end{remark}

\begin{remark}
Let $f\in \CC[x_1,\ldots,x_s]$ be a weighted homogeneous polynomial with respect to $w\in\NN^s$ and degree $d$ that defines an isolated hypersurface singularity. Consider any set $S$ of weighted homogeneous polynomials satisfying $d\leq \deg_w g$, for all $g\in S$. Then, proceeding exactly as in the proof of Theorem \ref{theorem-main}, it follows that $\CC[x_1,\ldots,x_s]/\langle f, S\rangle$ does not admit negative weighted derivations.
\end{remark}

\bibliographystyle{alpha}
\bibliography{References}

\end{document}